The Jurassic Zoo charges ​$12 for each adult admission and ​$3 for each child. The total bill for the 251 people from a school trip was ​$1455. How many adults and how many children went to the​ zoo? 

Answer by Boreal(15235)

x adults
251-x children
12x+3(251-x)=1455
12x+753-3x=1455
9x=702
x=78 adults ($936)
251-x=173 children ($519)